litnana17:

What is the radius of a cylinder if it has a volume of 769.3 cubic meters and a height of 5 meters? (Use 3.14 for π.)

darkknight:

volume of a cylinder \[\pi(r)^2h = 769.3\] We know what h is, now we need to find what what r is. Can you take it from here?

darkknight:

\[\pi(r)^25 = 769.3\]